OpenStudy (anonymous):

Hi I need help with this When a 2.75-kg fan, having blades 18.5 cm long, is turned off, its angular speed decreases uniformly from 10.0 rad/s to 6.30 rad/s in 5.00 s. (a) What is the magnitude of the angular acceleration of the fan? (b) Through what angle (in degrees) does it turn while it is slowing down during the 5.00 s? (c) If its angular acceleration does not change, how long after it is turned off does it take the fan to stop.

OpenStudy (anonymous):

i think it probolay c) in not that good at this stuff nut i try my best

OpenStudy (anonymous):

No you have to find a, b & c

OpenStudy (anonymous):

oh okey

OpenStudy (anonymous):

\[\alpha = \frac{ d \omega }{ dt }\] a) you can solve with this

OpenStudy (anonymous):

and for b) you need also \[\omega = \frac{ d ^{o} }{ dt }\]

OpenStudy (anonymous):

Thank you